OpenStudy (anonymous):

What's the first step in solving x/5 = 12/30? 12*x?

OpenStudy (anonymous):

Original: x/5 = 12/30 What you do to one side you have to do to the other, so: \[Step-1: 5 * (x/5) = (12/30) * 5\]To get rid of the 5 on the one side you have to apply the calculations to both sides, hence the multiplication to both sides.

OpenStudy (anonymous):

Wouldn't it be 12x=120?

OpenStudy (anonymous):

For 12 * x = 120, you can do a similar step, where you apply the calculation to both sides. \[Step-1: (12 * x)/12 =120/12\]This allows for you to solve for x.